OpenStudy (anonymous):

write an equation for the line containing (1,-11) and (2,-14)

OpenStudy (tkhunny):

Have you considered the Two-Point Form?

OpenStudy (anonymous):

whats that

OpenStudy (anonymous):

like the slope form

OpenStudy (anonymous):

yes find the slope

OpenStudy (anonymous):

and then apply it to y=mx+b

OpenStudy (anonymous):

y=mx+b y=-3x-8

OpenStudy (tkhunny):

Slope-Interecept is NOT directly applicable. You don't have an intercept.

OpenStudy (anonymous):

thats what i thought

OpenStudy (anonymous):

cuz its not looking for the slope

OpenStudy (tkhunny):

You can USE the slope-intercept form, but you don't KNOW the intercept, so you do have to go about it a little differently. (1,-11) and (2,-14) Find the slope (-14-(-11))/(2-1) = -3/1 = -1 Slope-Intercept Form y = -3x + b Substitute a point and solve for b. I'll use (1,-11) -11 = -3(1) + b -11 = -3 + b -8 = b Slope-Intercept Form y = -3x -8

OpenStudy (anonymous):

Thank you so so so so much! I've been waiting for a response for a long time. :)

OpenStudy (tkhunny):

Well, you had plenty of responses. I just finally did the whole problem, since the conversation had wandered around a bit. Always find hints, no matter the response.